# Information Re-Organization Improves Reasoning in Large Language Models

### v1

@@ 3.3 Reasoning With Re-Organized Information
please answer the question based on the
Context.
This prompt can be combined with existing methods such as Chain-of-Thought (CoT) Wei et al. (2022) by appending “Let’s think step by step.” to the question to further improve the accuracy of reasoning. Then we use a regex expression to extract answer $\mathcal{A}$ from the response.
4 Experimental
4.1 Tasks and Datasets
To verify the effectiveness of our information re-organization method, we conduct experiments across a range of contextually aware multi-hop reasoning tasks and datasets, including claim verification Guo et al. (2021), question answering Perez et al. (2020), and reading comprehension Dua et al. (2019). The detailed dataset information we used is displayed in Table 1.
Task Dataset Split Pairs
Claim Verification FEVEROUS Dev 2959
HOVER Dev 4000
SCIFACT Test 212
Question Answering 2WMHopQA Dev 500
StrategyQA Dev 229
MusiQue Dev 2417
Reading Comprehension HotpotQA Dev 500
WIKIHOP Dev 500
Claim Verification
The task involves assessing a given claim against a set of evidence documents to determine whether they support or refute the claim Guo et al. (2021). We consider HOVER Jiang et al. (2020) and FEVEROUS Aly et al. (2021), which comprise complex claims that necessitate multi-hop reasoning for verification. Additionally, we also take into account the SCIFACT Wadden et al. (2022) dataset, notable for its inclusion of scientific claims.
LLMs Methods HOVER (2-hop) HOVER (3-hop) HOVER (4-hop) FEVEROUS SCIFACT
LLAMA2-70B Standard 49.41 48.35 47.82 63.39 60.70
InfoRE 51.66 50.30 49.45 67.01 63.39
$\uparrow 2.25$ $\uparrow 1.95$ $\uparrow 1.63$ $\uparrow 3.62$ $\uparrow 2.69$
CoT 50.02 48.76 48.01 64.53 61.24
InfoRE + CoT 52.40 51.21 50.07 67.96 64.24
$\uparrow 2.38$ $\uparrow 2.45$ $\uparrow 2.06$ $\uparrow 3.43$ $\uparrow 3.00$
GPT-3.5 Standard 66.77 66.10 63.91 92.00 77.42
InfoRE 69.10 67.03 64.39 94.00 80.65
$\uparrow 2.33$ $\uparrow 0.93$ $\uparrow 0.48$ $\uparrow 2.00$ $\uparrow 3.23$
CoT 68.66 66.78 64.16 92.53 75.27
InfoRE + CoT 70.24 67.86 65.16 94.32 81.72
$\uparrow 1.58$ $\uparrow 1.08$ $\uparrow 1.00$ $\uparrow 1.79$ $\uparrow 6.45$
GPT-4 Standard 72.40 71.02 70.06 92.33 91.40
InfoRE 75.42 73.95 71.22 94.67 93.55
$\uparrow 3.02$ $\uparrow 2.93$ $\uparrow 1.16$ $\uparrow 2.34$ $\uparrow 2.15$
CoT 73.82 72.07 70.68 92.67 92.47
InfoRE + CoT 75.87 74.26 72.08 95.64 94.77
$\uparrow 2.05$ $\uparrow 2.19$ $\uparrow 1.40$ $\uparrow 2.97$ $\uparrow 2.30$
Question Answering
For this task, we consider the following datasets: 2WikiMultiHopQA Ho et al. (2020), StrategyQA Geva et al. (2021), MuSiQue Trivedi et al. (2021), and HotpotQA Yang et al. (2018). To answer the questions in these datasets requires not only multi-hop reasoning but also cross-document analysis.
Reading Comprehension
Machine reading comprehension task requires a model to process documents and select an answer from the provided candidates to a question about the content Dua et al. (2019). We primarily consider WIKIHOP Welbl et al. (2017) in the task, which necessitates multi-hop reasoning to derive the final answer. Additionally, HotpotQA dataset Yang et al. (2018) is also frequently considered as part of the question answering domain.
4.2 Language Models and Baselines
Models
In our paper, we employ pre-trained large language models such as Llama2-70B Touvron and et al. (2023), GPT-3.5 (text-davinci-003) Ouyang et al. (2022) and GPT-4 OpenAI (2023) as the foundational models. We configure all models with top_p parameter as $1.0$ and temperature as $0.0$.
Baselines
1) Standard. 2) CoT. The Standard approach is a method that directly reasons with the original textual context. The Chain of Thought (CoT) Wei et al. (2022) method involves augmenting standard reasoning methods by adding a step-by-step thought process. In our paper, we conduct the CoT strategy by appending the sentence “Let’s think step by step.” at the end of the question.
In this paper, we conduct experiments, including both the baseline methods and our InfoRE in a zero-shot setting to counteract the potential randomness associated with demonstrations in a few-shot setting. We also design an answer-format instruction within the prompts for various tasks to standardize the structure of the final answer, thereby enhancing the precision of answer extraction. Moreover, all results reported in the paper use only the reorganized contextual information to reason. Comprehensive details about prompts and answer-format instruction are available in Appendix B.
4.3 Evaluation Metrics
We use F1 to measure the results on all the datasets we used in our paper following previous methods Pan et al. (2023); Radhakrishnan et al. (2023); Aly et al. (2021). We run the official evaluation scripts of each dataset to get the output scores.
LLMs Methods 2WikiMultiHopQA MuSiQue StrategyQA HotpotQA WIKIHOP
LLAMA2-70B Standard 52.56 49.55 51.23 66.07 40.32
InfoRE 56.83 52.25 54.86 69.73 42.02
$\uparrow 4.27$ $\uparrow 2.70$ $\uparrow 3.63$ $\uparrow 3.66$ $\uparrow 1.70$
CoT 52.99 52.90 56.80 66.80 41.07
InfoRE + CoT 57.67 56.07 59.36 70.32 42.77
$\uparrow 4.68$ $\uparrow 3.17$ $\uparrow 2.56$ $\uparrow 3.52$ $\uparrow 1.70$
GPT-3.5 Standard 58.25 55.01 59.39 73.30 48.92
InfoRE 63.80 57.38 62.78 76.53 51.13
$\uparrow 5.55$ $\uparrow 2.37$ $\uparrow 3.39$ $\uparrow 3.23$ $\uparrow 2.21$
CoT 59.37 57.05 67.51 73.90 49.65
InfoRE + CoT 64.51 59.87 69.87 77.20 52.05
$\uparrow 5.14$ $\uparrow 2.82$ $\uparrow 2.36$ $\uparrow 3.30$ $\uparrow 2.40$
GPT-4 Standard 72.69 62.65 68.32 79.33 55.46
InfoRE 76.27 66.02 70.46 82.81 57.66
$\uparrow 3.58$ $\uparrow 3.37$ $\uparrow 2.14$ $\uparrow 3.48$ $\uparrow 2.20$
CoT 74.08 64.36 68.50 80.66 56.02
InfoRE + CoT 78.43 68.82 70.88 83.75 58.32
$\uparrow 4.35$ $\uparrow 4.46$ $\uparrow 2.38$ $\uparrow 3.09$ $\uparrow 2.30$
Table 3: Zero-shot results on Question Answering and Reading Comprehension tasks.
5 Results
5.1 Claim Verification
Table 2 presents a comprehensive performance comparison between our InfoRE and existing zero-shot techniques. For the HOVER dataset, we segment it into the 2-hop, 3-hop, and 4-hop levels according to the hop number in the dataset following ProgramFC Pan et al. (2023). As depicted in Table 2, our InfoRE demonstrates significant improvements in the zero-shot claim verification task. The CoT Wei et al. (2022) approach offers a lightly increase of 0.62% in the HOVER 4-hop using GPT-4, which indicates its marginal utility in more complex reasoning scenarios. Yet, our InfoRE achieves 1.16% improvement on the HOVER 4-hop using GPT-4 showing remarkable performance on contextual aware understanding and reasoning. This improvement is further increased to 72.08% in combination with CoT, suggesting that the methods complement each other effectively.
In the case of Llama2-70B, the combined application of InfoRE and CoT yields a score of 52.40% on the 2-hop HOVER task, surpassing its CoT-only score of 50.02%. This pattern of improvement is consistent with GPT-3.5, where InfoRE and CoT together achieved the highest reported score of 70.24% on the same dataset. GPT-4 shows superior performance across all methods and datasets. With the Standard approach, it reaches 70.06% accuracy in the 4-hop challenge, suggesting an inherent advanced reasoning ability. This performance is mirrored in the specialized benchmarks, where GPT-4 with InfoRE attains near-perfect accuracy on FEVEROUS (94.67%) and very high accuracy on SCIFACT (93.55%). These figures solidify the notion that the InfoRE indeed enhances the reasoning capabilities of LLMs.
5.2 Question Answer
Different from claim verification datasets, this task involves using multiple documents as context, presenting a challenge in cross-document reasoning. Intuitively, when the reasoning process involves multiple documents, information re-organization can effectively merge the information from different documents and uncover logical relationships that are not apparent in plain text. Performance in Table 3 verifies the intuitive. We can see that after applying information re-organization, all the results have a significant improvement. GPT-4 outperforms the other models, with the highest F1 score of 76.27% when employing InfoRE on 2WikiMultiHopQA, and results on GPT-3.5 also show a notable increase in performance with InfoRE, achieving a 64.51% F1 score, surpassing its Standard and CoT results. The StrategyQA dataset, known for requiring strategic reasoning, GPT-4 with InfoRE reaching an F1 of 70.46%. These scores are the highest across all models and methods, with GPT3.5 also showing a solid increase with InfoRE at 62.78% F1. The HotpotQA results further affirm the superiority of GPT-4 using InfoRE, which achieves an F1 of 83.75%. This is consistent with the performance improvements observed with GPT3.5 (text-davinci-003) and Llama2-70B when applying InfoRE.
5.3 Machine Reading Comprehension
Different from conventional question-answering tasks, this task includes a set of candidate answers, and the goal of the model is to identify the correct answer from these options. This not only requires LLMs to deeply understand the contextual content but also to accurately match the answers with the available candidates. Moreover, understanding candidate answers is crucial because they are sometimes disruptive to correct answers, consequently increasing the difficulty of the reasoning task. Table 3 illustrates performance on the WIKIHOP and HotPotQA datasets. Our InfoRE consistently shows improvements across all datasets. Specifically, in the WIKIHOP dataset, GPT-3.5 with InfoRE outperforms other methods with an F1 of 51.13%. This improvement further verifies the effectiveness of our method.
6 Analysis
To verify that information re-organization enables LLMs to improve reasoning ability, we conduct a series of analysis experiments on the question answer and claim verification tasks. We first evaluate the quality of re-organized information with GPT-4 (gpt-4-32k) OpenAI (2023) in Section 6.1 Then we analyze the impact of the quality of re-organized information on reasoning in Section 6.2 and conduct a detailed error analysis of the predictions in Section 6.3.
6.1 Quality of Re-Organized Information
To assess the quality of the re-organized information, we perform a quantitative evaluation of the re-organized information with GPT-4 (gpt-4-32k) on the 2WikiMultiHopQA dataset. Specifically, we select 100 samples from the dataset, GPT-4 (gpt-4-32k) is asked to rank re-organized information produced by GPT-3.5 (text-davinci-003) Brown et al. (2020) and GPT-4, as well as original context following criteria: (1) Depth: The information present multiple relationships of a topic, offering insightful perspectives or in-depth understanding of the subject. (2) Clarity: Information is clear and precise, making it easy to understand without ambiguity, and contributes directly to the overall message. All of the information is ranked 1, 2, and 3 with 3, 2, and 1 scores, respectively. Finally, we get a weighted average score for each information to measure the overall quality.
Methods Depth
1st 2nd 3rd Avg R.
Original 0.22 0.36 0.42 1.80
GPT-3.5 0.34 0.35 0.31 2.03
GPT-4 0.44 0.29 0.27 2.17
Methods Clarity
1st 2nd 3rd Avg R.
Original 0.25 0.35 0.40 1.85
GPT-3.5 0.35 0.33 0.32 2.03
GPT-4 0.40 0.32 0.28 2.12
Results are shown in Table 4. The results demonstrate that the re-organized information outperforms the original textual context in terms of depth and clarity, which justifies the motivation of our paper. In addition, the re-organized information from GPT-4 outperforms GPT-3.5, which proves in another way that GPT-4 is indeed more capable than GPT-3.5. The evaluation results further validate the effectiveness of our method. Furthermore, the 20.56% improvement in depth is more obvious than the 14.59% improvement in clarity, which indicates that the re-organization of the information has been particularly effective in enhancing the logical relationships of the information. The improvement in clarity suggests that the information is now presented in a more direct and streamlined manner, making it easier for LLMs to grasp the essential points without wading through unnecessary details.
6.2 Effect of Re-organized Information Quality
Methods FEVEROUS 2WikiMultiHopQA
GPT-3.5 (†)
  Standard 92.00 58.25
  InfoRE 94.00 63.80
  InfoRE * 94.52 65.86
GPT-4 (*)
  Standard 92.33 72.69
  InfoRE 94.67 76.27
  InfoRE 94.32 74.82
The cross-validation results, shown in Table 5, indicate that the use of either GPT-3.5 or GPT-4 for information re-organization leads to a marked improvement in the performance of LLMs on reasoning tasks. Additionally, it is observed that using GPT-4 for information re-organization while reasoning with the GPT-3.5 results in a further increase in reason results by 0.52% and 2.06% on FEVEROUS and 2WikiMultiHopQA datasets, respectively. This implies that enhancing the quality of re-organized information can improve the performance of LLMs, pointing to a promising direction for future research in refining information synthesis and re-organization. Conversely, when GPT-3.5 is employed for information re-organization in conjunction with reasoning using the GPT-4 model, there is a decrease in reason ability by 0.35% and 1.45% respectively. In addition, the magnitude of the decrease is lower than the increase, which implies that our re-organization strategy may have a greater impact on models with weaker reasoning abilities. This finding aligns with our understanding that models with inherently weaker reasoning abilities tend to rely more heavily on external strategy. In contrast, models with stronger inherent capabilities can use their strengths to mitigate the adverse effects.
6.3 Error Analysis
To better comprehend where the errors in our InfoRE methodology come from and where they are fixed, we annotate 100 wrong predictions made by both InfoRE and Standard methods with GPT-3.5 (text-davinci-003) on 2WikiMultiHopQA dataset. We categorize the errors into $4$ classes:
  • Contextual Misunderstanding This happens when the model fails to interpret or connect multiple pieces of information from different parts of the documents. Multi-hop reasoning requires synthesizing information from various segments, and recognizing logical relations, and any misunderstanding can lead to incorrect conclusion.
  • Factual Error The model may provide an answer that is factually incorrect or not supported by the given documents. This is often due to the model’s reliance on its training data, which may not always align with the specific facts in the context.
  • Mathematical Error The error occurs when math calculations are involved in deriving the final answer.
  • Unanswerable Question It’s a specific type of error or limitation in dataset design, where the context does not contain enough information to provide a valid answer to the posed question.
When engaging in reasoning with large language models, all four error categories are present. As shown in Figure 3, there are 6% unanswerable question errors in the dataset, and more than 90% errors occur in the reasoning in the baseline method. Among the four types of error, contextual misunderstanding is the primary source of errors in the baseline, highlighting the importance of an in-depth understanding of context in solving reasoning tasks with large language models. This finding is consistent with our motivation presented in the introduction section. Moreover, comparing the results of errors between our method and the baseline method, our method mainly corrects 13% of errors coming from the baseline method, most of the corrected errors are contextual misunderstanding errors. This further indicates that our InfoRE method assists Large Language Models (LLMs) in understanding context, signifying the necessity and effectiveness of conducting information re-organization before directly addressing the original question.
Figure 3: Error Analysis of InfoRE on 2WikiMultiHopQA against Standard baseline method. The first four rectangles are error categories, while “Corrected” on the far right denotes the percentage of errors originally made by the baseline method that our method InfoRE has successfully corrected.
7 Conclusion
In this paper, we propose an information re-organization method to improve the reasoning ability of large language models (LLMs). Compared with previous approaches primarily focus on improving the quality of intermediate steps, our method emphasizes uncovering the logical relationships and multi-hop connections by information re-organization. By re-organizing the way information is presented, we aim to enable LLMs to explicitly perceive the logical relationships and multi-hop connection of concepts in the given context. This facilitates a deeper integration and understanding of context, leading to more robust reasoning results. To verify the effectiveness of our method, we conduct experiments using various LLMs across a range of contextually aware multi-hop reasoning tasks. The experiment results demonstrate the potential of our method to improve the reasoning ability of LLMs.
Appendix A Prompt for Obtaining MindMap
The specific prompt to obtain MindMap is shown in Table  6. In the prompt, the content within the brackets should be replaced with specific example content.
Prompt Content for Obtaining MindMap
Given a claim and corresponding evidence, please summarize the evidence as a mind map according to the claim. The output must be in a strict JSON format: {“mind_map”: “mind_map”}. CLAIM: [CLAIM] EVIDENCE: [EVIDENCE] .
Table 6: Specific prompt for obtaining MindMap. During execution, “[EVIDENCE]” needs to be replaced with the specific document, and “[CLAIM]” is replaced with the specific question.
Appendix B Prompt for Multi-hop Reason
The specific prompts for Standard, chain-of-thought (CoT), our InfoRE, and InfoRE + CoT to reason are shown in Table  7. In the prompt, the content within the brackets [] should be replaced with specific example content. Specifically, in the claim verification task, [CLAIM] corresponds to the claim of the specific example, and [EVIDENCE] corresponds to the evidence of the specific example. In the question-answering task, [CLAIM] corresponds to the question of the specific example, and [EVIDENCE] corresponds to the context document. [MindMap] corresponds to the MindMap of the specific example.
@@ Appendix B Prompt for Multi-hop Reason
Table 7: Specific prompts and answer format instructions of Standard, CoT, InforRE, and InfoRE + CoT in our paper. During execution, “[EVIDENCE]” needs to be replaced with the specific document, “[CLAIM]” is replaced with the specific question, and “[MindMap]” is replaced with the specific MindMap.

### v2

@@ Appendix C Prompt for Multi-hop Reason
please answer the question based on the documents.
Answer:
The specific prompts for Standard, chain-of-thought (CoT), our InfoRE, and InfoRE + CoT to reason are shown in Table  7. In the prompt, the content within the brackets [] should be replaced with specific example content.
Table 7: Specific prompts and answer format instructions of Standard, CoT, InforRE, and InfoRE + CoT in our paper. During execution, “[EVIDENCE]" needs to be replaced with the specific document, “[CLAIM]" is replaced with the specific question, and “[MindMap]" is replaced with the specific MindMap.
@@
Appendix D Detailed Dataset Information
In our experiments, due to the resource limitations of large language models, we sample a portion from each dataset, following previous methods [6, 16].
Tasks Dataset Pairs
train test
Claim Verification FEVEROUS 2000 2959
HOVER 2000 4000
SCIFACT 200 212
Question Answering 2WikiMultiHopQA 2000 500
StrategyQA 1000 229
MusiQue 2000 2417
Reading Comprehension HotpotQA 2000 500
WIKIHOP 2000 500
Claim Verification
The task involves assessing a given claim against a set of evidence documents to determine whether they support or refute the claim [12]. We consider HOVER [24] and FEVEROUS [25], which comprise complex claims that necessitate multi-hop reasoning for verification. Besides, we also take into account the SCIFACT [26] dataset, notable for its inclusion of scientific claims.
Question Answering
For this task, we consider the following datasets: 2WikiMultiHopQA [27], StrategyQA [28], MuSiQue [29], and HotpotQA [30]. To answer the questions in these datasets requires not only multi-hop reasoning but also cross-document analysis.
Reading Comprehension
Machine reading comprehension task requires a model to process documents and select an answer from the provided candidates to a question about the content [14]. We primarily consider WIKIHOP [31] in the task, which necessitates multi-hop reasoning to derive the final answer. Additionally, HotpotQA dataset [30] is also frequently considered as part of the question answering domain.
Appendix E Limitations
We propose an information re-organization approach to improve the reasoning of large language models, which performs well on some context-aware reasoning tasks but still has some limitations. Firstly, the structures of information re-organization are limited. Generally, there are multiple structures for information reorganization, such as tables, timelines, etc. Next, we will extend the re-organization structure to include more types. Secondly, the re-organization process relies on large language models. If we can implement this re-organization using smaller language models, our method will become more generalizable. This is another direction we need to focus on in the future.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: Abstract, Introduction
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: Limitations in Appendix
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory Assumptions and Proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: The paper does not include theoretical results
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental Result Reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: Experimental
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: Supplementary material
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental Setting/Details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: Experimental
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment Statistical Significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [No]
  34. Justification: Error bars are not reported because it would be too computationally expensive.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments Compute Resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: Experimental
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code Of Ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: Supplementary material
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader Impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: Conclusion
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: The paper poses no such risks.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: Introduction, Related Work, Methodology, Experimental, Results and Analysis.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New Assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: The paper does not involve the release of new assets.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and Research with Human Subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.